# Ammann Bars for Octagonal Tilings

### v4

@@ Abstract
Ammann bars are formed by segments (decorations) on the tiles of a tiling such that forming straight lines with them while tiling forces non-periodicity. Only a few cases are known, starting with Robert Ammann’s observations on Penrose tilings, but there is no general explanation or construction. In this article we propose a general method for cut and project tilings based on the notion of subperiods and we illustrate it with an aperiodic set of 36 decorated prototiles related to what we called Cyrenaic tilings.
@@ 1 Introduction
Shortly after the famous Penrose tilings were introduced by Roger Penrose in 1974 [23, 24, 25] and popularized by Martin Gardner in 1977 [15], amateur mathematician Robert Ammann [29] found particularly interesting decorations of the tiles (Figure 1): if one draws segments in the same way on all congruent tiles333taking into account the original markings on the sides that define the assembly rules. then on any valid tiling all those segments compose straight lines, going in five different directions. Conversely if one follows the assembly rule consisting of prolonging every segment on the tiles into a straight line then the obtained tiling is indeed a Penrose tiling. Those lines are called Ammann bars and the corresponding matching rule is locally equivalent to the ones given by Penrose using arrows on the sides or alternative decorations.
@@
Penrose tilings have many interesting properties and can be generated in several ways. The cut and project method [^4] Terms in italic which are not defined in the introduction are defined formally in further sections. The introduction is meant to give a general idea of the article. follows their algebraic study by de Bruijn in 1981 [10]. Beenker soon proposed a whole family of tilings based on it [8], including the Ammann-Beenker tilings that Ammann found independently. A cut and project tiling can be seen as a digitization of a two-dimensional plane in a $n$-dimensional Euclidean space ($n>2$), and we will talk about $n\to 2$ tilings in that sense. When the slope of the plane does not contain any rational line, the tiling is non-periodic. This is the case for Penrose tilings for instance, so the set of tiles defining them is aperiodic: one can tile the plane with its tiles but only non-periodically. The first aperiodic tileset was found by Berger, thus proving the undecidability of the Domino Problem [9] and relating tilings to logic. Since then, relatively few others were exhibited: many non-periodic tilings exist (even infinitely many using the cut and project method), but we usually do not have a corresponding aperiodic tileset.
@@
Links were made between such tilings and quasicrystals [28, 32], that is crystals whose diffraction pattern is not periodic but still ordered, with rotational symmetries. The study of local rules, i.e. constraints on the way tiles can fit together in finite patterns, can help modeling the long range aperiodic order of quasicrystals. For instance, Penrose tilings are defined by their 1-atlas [12], which is a small number of small patterns: any and all tilings containing only those patterns (of the given size) are Penrose tilings. Alternately, they can also be defined by their Ammann local rules, as stated in the first paragraph. On the contrary, it was proven [11] that Ammann-Beenker tilings, also known as 8-fold tilings, do not have weak local rules (or “colored” local rules), i.e. no finite set of patterns (r-atlas) is enough to characterize them. Socolar found sort of Ammann bars for them [30], but they extend outside the boundary of the tiles, thus do not fit the framework considered here.
@@
In addition to finding a suited tileset, one of the difficulties lies in proving that this tileset, with only nearest-neighbor rules, can enforce projection tilings with arbitrary slopes. Goodman-Strauss [16] and then Fernique and Ollinger [13] designed constructions in the case of substitutive tilings: they showed that, under specific conditions, substitutive tilings have colored nearest-neighbor matching rules. In other words, substitutions can be forced by colored local rules. In the case of cut and project tilings, this is a corollary of a construction presented later by Fernique and Sablik [14] and based on works by Aubrun and Sablik [1], and Hochman [19]. This more general construction yields weak local rules for every computable slope, yet it is quite heavy, since it simulates Turing machines. The main result presented here is actually a corollary of Fernique and Sablick’s, but the construction is completely different and much lighter because specific to a narrower class of tilings, for (a priori) non-substitutive tilings, and additionally we exhibit more information on the structure of these tilings in the process. Namely, we provide sufficient conditions for some cut and project tilings to have Ammann bars.
@@
Grünbaum and Shephard [17] detail the properties of Ammann bars in the case of Penrose tilings and their close relation to the Fibonacci word, which inspired Stehling to find an equivalent in the 3-dimensional space [31]. They also present two tilesets by Ammann with Ammann bars (A2 and A3) but these are substitutive and not cut and project tilings. Generally speaking, we do not know much about Ammann bars and for now each family of aperiodic tilings has to be treated on a case-by-case basis, as Lück did in [22] with the examples of Ammann bars introduced by Socolar [30] and Ingalls [20]. Yet they can reveal quite useful to study the structure of tilings, and were used by Porrier and Blondin Massé [26] to solve a combinatorial optimization problem on graphs defined by Penrose tilings.
@@
When it comes to $4\to 2$ tilings (digitizations [^5] The formal notion is planar, as defined in Section 2. of 2-dimensional planes in $\mathbb{R}^{4}$) and a few others like Penrose, which are $5\to 2$ tilings, the existence of weak local rules can be expressed in terms of subperiods, which are particular vectors of the slope [4, 6]. Careful observation of Penrose tilings from this angle shows that Ammann bars have the same directions as subperiods: there are two subperiods in each direction, one being $\varphi$ times longer than the other. Additionally, the lengths of the “integer versions” of subperiods are closely related to the distances between two consecutive Ammann bars in a given direction, as can be seen in Figure 1 (more details are given in Appendix A). Though interesting, this special case is too particular to hope for a generalization from it alone. Nonetheless, we think that Ammann bars are related to subperiods.
@@
In any case, it seems some conditions of alignment play a part in the existence of Ammann bars. This led us to introduce the notion of fine projection (Def. 1 p. 1) on a slope. We propose a constructive method – the FP-method (fine projection method), to find Ammann bars for $4\to 2$ tilings which are characterized by subperiods and for which we can find a fine projection. The FP-method outputs a finite set of tiles, which we call an FP-tileset. We prove the following result:
@@
We initially searched for examples of slopes based on two criteria: the slope has to be quadratic to meet the algebraic constraints (being characterized by subperiods), and we wanted “short” subperiods to facilitate observations on drawings. The method involves a trial-and-error component: for each slope, we have to test whether it is possible to find a fine projection or not. Our first try was on Golden Octagonal tilings but their slope admits no fine projection. Hence we randomly generated potential candidates, found a few ones having a fine projection, and proved the above result on a particular example which we presented in the LATIN 2022 conference [27]. It consists of $4\to 2$ tilings defined by a given slope based on the irrationality of $\sqrt{3}$, that we called Cyrenaic tilings in reference to Theodorus of Cyrene who proved $\sqrt{3}$ to be irrational. In this case, the FP-method yields the set of decorated tiles depicted in Figure 2. Those tiles give Ammann bars to Cyrenaic tilings.
@@
We reuse this example here to illustrate the general case. We then found many examples of $4\to 2$ tilings characterized by their subperiods and admitting a fine projection, and we provide a program to generate the corresponding decorated tilesets for any slope meeting these conditions. Note that there might be several fine projections for a given slope, hence different possible decorated tilesets depending on the choice we make at this step.
The case of Penrose indicates that our construction could (and should) be adapted in order to work for $5\rightarrow 2$ tilings, or general cut and project ($n\to d$) tilings. In particular, for Penrose the lines are shifted and the number of lines is reduced compared with our method, so that only two decorated tiles are needed. Besides, in each direction the distance between two consecutive lines can take only two values, and the sequence of intervals is Sturmian and substitutive. In the case of Cyrenaic tilings, it is an open question whether the bi-infinite word defined by each sequence of intervals between Ammann bars is substitutive. If so, maybe we could compose them after finding the substitution. As for $n\to 3$ tilings, the case of icosahedral Ammann rhombohedral tilings could be studied as a first example of “Ammann planes” (instead of lines), since their slope is characterized by subperiods [5].
More generally, we conjecture that the bi-infinite words defined by Ammann bars are either binary or ternary words, probably billiard words, in relation with the Three-Distance Theorem. Such questions are still open, even though we have clues as to how such properties could be explained. Lines could also be shifted as it is the case for Penrose tilings, instead of passing through vertices. An optimal shift (reducing the number of lines or tiles) would then have to be determined. Besides, it would be interesting to know whether the hypothesis of having a fine projection is necessary to have Ammann bars. A link to our SageMath code as well as some more technical explanations are given in Appendix B.
The paper is organized as follows. Section 2 introduces the settings, providing the necessary formal definitions, in particular local rules and subperiods. In Section 3 we present the FP-method to construct a set of decorated prototiles yielding Ammann bars. We rely on subperiods characterizing a slope as well as a fine projection, and prove that the obtained set is always finite. Finally, in Section 4 we show that Ammann segments on the tiles of an FP-tileset force the corresponding tilings to all have the same subperiods, thus proving Theorem 1.
2 Settings
@@ 2.2 Local rules
Local rules for tilings can be defined in several ways, which are not equivalent. Since we focus on cut and project tilings, we also define local rules for a slope.
@@ 2.3 Subperiods
An $i_{1},...,i_{n-3}$-subperiod of an $n\to 2$ tiling $T$ is a prime period of its $i_{1},...,i_{n-3}$-shadow, hence an integer vector in $\mathbb{R}^{3}$. By extension, we call subperiod of a slope $E$ any vector of $E$ which projects on a subperiod in a shadow of $T$. A subperiod is thus a vector of $E$ with 3 integer coordinates: those in positions $j\notin\{i_{1},...,i_{n-3}\}$. We say that a slope is determined or characterized by its subperiods if only finitely many slopes have the same subperiods (in the shadows). For instance, the slope of Ammann-Beenker tilings has four subperiods:
@@ 3 The FP-method
In this section, we present a construction to get Ammann bars for some $4\rightarrow 2$ tilings and we illustrate it with what we named Cyrenaic tilings.
@@ 3.1 Fine projections
In Subsection 2.1, we defined what is a valid projection for a slope $E$ and mentioned the classical case of the orthogonal projection. There are however other valid projections, and this will play a key role here. We will indeed define Ammann bars as lines directed by subperiods and it will be convenient for the projected $i$-th subperiod $\pi(p_{i})$ to be collinear with $\pi(e_{i})$, so that the image of a line directed by $p_{i}$ is still a line in the $i$-th shadow (Figure 6).
@@
Figure 7 illustrates the difference between two valid projections, one being fine but not the other, on the slope of Cyrenaic tilings which we present in the next subsection. With the fine projection, projected subperiods have the same directions as the sides of the tiles. This is why if segments on the tiles of a tiling $T$ are directed by $\pi(p_{i})$ then continuity of the lines in direction $i$ is preserved in the $i$-shadow of $T$, for any $i\in\{0,1,2,3\}$, as illustrated in Figure 6. Indeed, consider a line $L$ in direction $i$, then it is parallel to the sides of the tiles which disappear in the $i$-shadow of $T$. Now consider a tile $t_{0}$ which disappears in this shadow, containing a segment $[BC]\subset L$, and its neighbors $t_{-1}$ and $t_{1}$ containing segments $[AB],[CD]\subset L$. Taking the $i$-shadow corresponds to translating remaining tiles in direction $i$, hence by such a translation the endpoint of an Ammann segment is mapped to a point on the same line (namely the image of the other endpoint of the same segment). As a result, the images of points $B$ and $C$ are on the same line, so that points $A,B,C,D$ are still aligned.
@@ 3.2 Finding fine projections
Given a slope $E$ with subperiods $p_{0},\ldots,p_{3}$, we search for a fine projection $\pi$ as follows. We will define it by its $2\times 4$ matrix $A$, which must satisfy $Ae_{i}=\lambda_{i}Ap_{i}$ for $i=0,\ldots,3$, where $\Lambda:=(\lambda_{i})_{i=0,\ldots,3}$ is to be determined. With $M$ denoting the $4\times 4$ matrix whose $i$-th column is $e_{i}-\lambda_{i}p_{i}$, this rewrites $AM=0$. The $2$ rows of $A$ must thus be in the left kernel of $M$. Since the image of the facets in $E+[0,1]^{4}$ must cover $\mathbb{R}^{2}$, $A$ must have rank $2$. Hence the left kernel of $M$ must be of dimension at least $2$, that is, $M$ must have rank at most $2$. This is equivalent to saying that all the $3\times 3$ minors of $M$ must be zero. Each minor yields a polynomial equation in the $\lambda_{i}$’s. Any solution of the system formed by these equations yields a matrix $M$ whose left kernel can be computed. If the kernel is not empty, then any basis of it yields a suitable matrix $A$.
@@
To get an idea of the proportion of irrational slopes for which a fine projection exists, we ran some statisctics with the following procedure. For an integer parameter $k$, we randomly choose a quadratic polynomial $P$ with coefficients in $[-k,k]$ until we find an irreducible one, then two irrational numbers $x$ and $y$ in the quadratic field defined by $P$. We also randomly choose two vectors $u$ and $v$ in $\mathbb{Z}^{4}$ with coefficients in $[-k,k]$, and we replace the first coordinate in $u$ by $x$ and the second in $v$ by $y$. Finally, we apply the above method on the slope $E$ generated by $u$ and $v$. For $k=3$, out of 999 slopes, 222 were not characterized by subperiods and 116 admitted a fine projection. For $k=5$, out of 7015 slopes, 765 were not characterized by subperiods and 1000 admitted a fine projection [^6] We searched until we found 1000 fine projections..

### v5

@@ Abstract
Ammann bars are formed by segments (decorations) on the tiles of a tiling such that forming straight lines with them while tiling forces non-periodicity. Only a few cases are known, starting with Robert Ammann’s observations on Penrose tilings, but there is no general explanation or construction. In this article we propose a general method for cut and project tilings based on subperiods and we illustrate it with an aperiodic set of 36 decorated prototiles related to what we called Cyrenaic tilings.
@@ 1 Introduction
Shortly after the famous Penrose tilings were introduced by Roger Penrose in 1974 [23, 24, 25] and popularized by Martin Gardner in 1977 [15], amateur mathematician Robert Ammann [29] found particularly interesting decorations of the tiles (Figure 1): if one draws segments in the same way on all congruent tiles333taking into account the original markings on the sides that define the assembly rules. then on any valid tiling all those segments compose continuous straight lines, going in five different directions. Conversely if one follows the assembly rule consisting of prolonging every segment on the tiles into a straight line then the obtained tiling is indeed a Penrose tiling. Those lines are called Ammann bars and the corresponding matching rule is locally equivalent to the ones given by Penrose using arrows on the sides or alternative decorations.
@@
Penrose tilings have many interesting properties and can be generated in several ways. The cut and project method [^4] Terms in italics which are not defined in the introduction are defined formally in further sections. The introduction is meant to give a general idea of the article. follows their algebraic study by de Bruijn in 1981 [10]. Beenker soon proposed a whole family of tilings based on it [8], including the Ammann-Beenker tilings that Ammann found independently. A cut and project tiling can be seen as a discretization of a two-dimensional plane in an $n$-dimensional Euclidean space ($n>2$), and we will talk about $n\to 2$ tilings in that sense. When the slope of the plane does not contain any rational line, the tiling is non-periodic. This is the case for Penrose tilings for instance, so the set of tiles defining them is aperiodic: one can tile the plane with its tiles but only non-periodically. The first aperiodic tileset was found by Berger, thus proving the undecidability of the Domino Problem [9] and relating tilings to logic. Since then, relatively few others were exhibited: many non-periodic tilings exist (even infinitely many using the cut and project method), but we usually do not have a corresponding aperiodic tileset.
@@
Links were made between such tilings and quasicrystals [28, 32], that is crystals whose diffraction pattern is not periodic but still ordered, with rotational symmetries. The study of local rules, i.e. constraints on the way tiles can fit together in finite patterns, can help modeling the long range aperiodic order of quasicrystals. For instance, Penrose tilings are defined by their 1-atlas [12], which is a small number of small patterns: any and all tilings containing only those patterns (of the given size) are Penrose tilings. Alternately, they can also be defined by their Ammann local rules, as stated in the first paragraph. On the contrary, it was proven [11] that Ammann-Beenker tilings, also known as 8-fold tilings, do not have weak local rules (or “colored” local rules), i.e. no finite set of patterns (r-atlas) is enough to characterize them. Socolar found a variant of Ammann bars for them [30], but they extend outside the boundary of the tiles, thus do not fit the framework considered here.
@@
In addition to finding a suitable tileset, one of the difficulties lies in proving that this tileset, with only nearest-neighbor rules, can enforce projection tilings with arbitrary slopes. Goodman-Strauss [16] and then Fernique and Ollinger [13] designed constructions in the case of substitutive tilings: they showed that, under specific conditions, substitutive tilings have colored nearest-neighbor matching rules. In other words, substitutions can be forced by colored local rules. In the case of cut and project tilings, this is a corollary of a construction presented later by Fernique and Sablik [14] and based on works by Aubrun and Sablik [1], and Hochman [19]. This more general construction yields weak local rules for every computable slope, yet it is quite heavy, since it simulates Turing machines. The main result presented here is actually a corollary of Fernique and Sablik’s, but the construction is completely different and much lighter because it is specific to a narrower class of tilings, for (a priori) non-substitutive tilings, and additionally we exhibit more information on the structure of these tilings in the process. Specifically, we provide sufficient conditions for some cut and project tilings to have Ammann bars.
@@
Grünbaum and Shephard [17] detail the properties of Ammann bars in the case of Penrose tilings and their close relation to the Fibonacci word, which inspired Stehling to find an equivalent in 3-dimensional space [31]. They also present two tilesets by Ammann with Ammann bars (A2 and A3) but these are substitutive and not cut and project tilings. Generally speaking, many questions remain around Ammann bars and for now each family of aperiodic tilings has to be treated on a case-by-case basis, as Lück did in [22] with the examples of Ammann bars introduced by Socolar [30] and Ingalls [20]. Yet they can prove quite useful in the study of the structure of tilings, and were used by Porrier and Blondin Massé [26] to solve a combinatorial optimization problem on graphs defined by Penrose tilings.
@@
When it comes to $4\to 2$ tilings (discretizations [^5] A discretization of a 2-plane $P$ is, roughly speaking, a set of unit facets of $\mathbb{Z}^{4}$ which approximate the continuous plane $P$. The formal notion is planar, as defined in Section 2. of 2-dimensional planes in $\mathbb{R}^{4}$) and a few others like Penrose, which are $5\to 2$ tilings, the existence of weak local rules can be expressed in terms of subperiods, which are particular vectors of the slope [4, 6]. Careful examination of Penrose tilings from this angle shows that Ammann bars have the same directions as subperiods: there are two subperiods in each direction, one being $\varphi$ times longer than the other. Additionally, the lengths of the “integer versions” of subperiods are closely related to the distances between two consecutive Ammann bars in a given direction, as can be seen in Figure 1 (more details are given in Appendix A). Though interesting, this special case is too particular to hope for a generalization from it alone. Nonetheless, we think that Ammann bars are related to subperiods.
@@
In any case, it seems some conditions of alignment play a part in the existence of Ammann bars. This led us to introduce the notion of fine projection (Def. 1 p. 1) on a slope. We propose a constructive method – the FP-method (fine projection method) – to find Ammann bars for $4\to 2$ tilings which are characterized by subperiods and for which we can find a fine projection. The FP-method outputs a finite set of tiles, which we call an FP-tileset. We prove the following result:
@@
We initially searched for examples of slopes based on two criteria: the slope has to be quadratic to meet the algebraic constraints (being characterized by subperiods), and we wanted “short” subperiods to enable human-eye observation of the corresponding drawings. The method involves a trial-and-error component: for each slope, we have to test whether it is possible to find a fine projection or not. Our first try was on Golden Octagonal tilings but their slope admits no fine projection. Hence we randomly generated potential candidates, found a few having a fine projection, and proved the above result on a particular example which we presented in the LATIN 2022 conference [27]. It consists of $4\to 2$ tilings defined by a given slope based on the irrationality of $\sqrt{3}$, that we called Cyrenaic tilings in reference to Theodorus of Cyrene who proved $\sqrt{3}$ to be irrational. In this case, the FP-method yields the set of decorated tiles depicted in Figure 2. Those tiles give Ammann bars to Cyrenaic tilings.
@@
We use this example here to illustrate the general case. We then found many examples of $4\to 2$ tilings characterized by their subperiods and admitting a fine projection, and we provide a program to generate the corresponding decorated tilesets for any slope meeting these conditions. Note that there might be several fine projections for a given slope, hence different possible decorated tilesets depending on the choice we make at this step.
The case of Penrose indicates that our construction could (and should) be adapted in order to work for $5\rightarrow 2$ tilings, or general cut and project ($n\to d$) tilings. In particular, for Penrose the lines are shifted and the number of lines is reduced compared with our method, so that only two decorated tiles are needed. Besides, in each direction the distance between two consecutive lines can take only two values, and the sequence of intervals is Sturmian and substitutive. In the case of Cyrenaic tilings, it is an open question whether the bi-infinite word defined by each sequence of intervals between Ammann bars is substitutive. If so, it may be that they could be composed according to the substitution. As for $n\to 3$ tilings, the case of icosahedral Ammann rhombohedral tilings could be studied as a first example of “Ammann planes” (instead of lines), since their slope is characterized by subperiods [5].
More generally, we conjecture that the bi-infinite words defined by Ammann bars are either binary or ternary words, probably billiard words, in relation with the Three-Distance Theorem. Such questions are still open, although there are clues as to how such properties could be explained. Lines could also be shifted as in the case of Penrose tilings, instead of passing through vertices. An optimal shift (reducing the number of lines or tiles) would then have to be determined. Besides, it would be interesting to know whether the hypothesis of having a fine projection is necessary to have Ammann bars. A link to our SageMath code as well as some more technical explanations are given in Appendix B.
The paper is organized as follows. Section 2 introduces the setting, providing the necessary formal definitions, in particular local rules and subperiods. In Section 3 we present the FP-method to construct a set of decorated prototiles yielding Ammann bars. We rely on subperiods characterizing a slope as well as a fine projection, to prove that the obtained set is always finite. Finally, in Section 4 we show that the Ammann segments on the tiles of an FP-tileset force the corresponding tilings to all have the same subperiods, thus proving Theorem 1.
2 setting
@@ 2.2 Local rules
Local rules for tilings can be defined in several inequivalent ways. Since we focus on cut and project tilings, we also define local rules for a slope.
@@ 2.3 Subperiods
An $i_{1},...,i_{n-3}$-subperiod of an $n\to 2$ tiling $T$ is a prime period of its $i_{1},...,i_{n-3}$-shadow, hence an integer vector in $\mathbb{R}^{3}$. By extension, we call subperiod of a slope $E$ any vector of $E$ which projects on a subperiod in a shadow of $T$. A subperiod is thus a vector of $E$ with 3 integer coordinates: those in positions $j\notin\{i_{1},...,i_{n-3}\}$. We say that a slope is determined or characterized by its subperiods if only finitely many slopes have the same subperiods (in their shadows). For instance, the slope of Ammann-Beenker tilings has four subperiods:
@@ 3 The FP-method
In this section, we present a construction to get Ammann bars for some $4\rightarrow 2$ tilings and we illustrate it with tilings that we call Cyrenaic tilings.
@@ 3.1 Fine projections
In Subsection 2.1, we defined the valid projections for a slope $E$ and mentioned the classical case of the orthogonal projection. There are however other valid projections, and this will play a key role here. Indeed, we will define Ammann bars as lines directed by subperiods and it will be convenient for the projected $i$-th subperiod $\pi(p_{i})$ to be collinear with $\pi(e_{i})$, so that the image of a line directed by $p_{i}$ is still a line in the $i$-th shadow (Figure 6).
@@
Figure 7 illustrates the difference between two valid projections, one being fine but not the other, on the slope of Cyrenaic tilings which we present in the next subsection. With the fine projection, projected subperiods have the same directions as the sides of the tiles. This is why, if segments on the tiles of a tiling $T$ are directed by $\pi(p_{i})$, then continuity of the lines in direction $i$ is preserved in the $i$-shadow of $T$, for any $i\in\{0,1,2,3\}$, as illustrated in Figure 6. Indeed, consider a line $L$ in direction $i$, then it is parallel to the sides of the tiles which disappear in the $i$-shadow of $T$. Now consider a tile $t_{0}$ which disappears in this shadow, containing a segment $[BC]\subset L$, and its neighbors $t_{-1}$ and $t_{1}$ containing segments $[AB],[CD]\subset L$. Taking the $i$-shadow corresponds to translating remaining tiles in direction $i$, hence by such a translation the endpoint of an Ammann segment is mapped to a point on the same line (namely the image of the other endpoint of the same segment). As a result, the images of points $B$ and $C$ are on the same line, so that points $A,B,C,D$ are still aligned.
@@ 3.2 Finding fine projections
Given a slope $E$ with subperiods $p_{0},\ldots,p_{3}$, we search for a fine projection $\pi$ as follows. We will define it by its $2\times 4$ matrix $A$, which must satisfy $Ae_{i}=\lambda_{i}Ap_{i}$ for $i=0,\ldots,3$, where $\Lambda:=(\lambda_{i})_{i=0,\ldots,3}$ is to be determined. With $M$ denoting the $4\times 4$ matrix whose $i$-th column is $e_{i}-\lambda_{i}p_{i}$, this can be rewritten as $AM=0$. The $2$ rows of $A$ must thus be in the left kernel of $M$. Since the image of the facets in $E+[0,1]^{4}$ must cover $\mathbb{R}^{2}$, $A$ must have rank $2$. Hence the left kernel of $M$ must be of dimension at least $2$, that is, $M$ must have rank at most $2$. This is equivalent to saying that all the $3\times 3$ minors of $M$ must be zero. Each minor yields a polynomial equation in the $\lambda_{i}$’s. Any solution of the system formed by these equations yields a matrix $M$ whose left kernel can be computed. If the kernel is not empty, then any basis of it yields a suitable matrix $A$.
@@
To get an idea of the proportion of irrational slopes for which a fine projection exists, we ran some statistics with the following procedure. For an integer parameter $k$, we randomly choose a quadratic polynomial $P$ with coefficients in $[-k,k]$ until we find an irreducible one, then two irrational numbers $x$ and $y$ in the quadratic field defined by $P$. We also randomly choose two vectors $u$ and $v$ in $\mathbb{Z}^{4}$ with coefficients in $[-k,k]$, and we replace the first coordinate in $u$ by $x$ and the second in $v$ by $y$. Finally, we apply the above method on the slope $E$ generated by $u$ and $v$. For $k=3$, out of 999 slopes, 222 were not characterized by subperiods and 116 admitted a fine projection. For $k=5$, out of 7015 slopes, 765 were not characterized by subperiods and 1000 admitted a fine projection [^6] We searched until we found 1000 fine projections..